if a man mixes 12 lbs. of peanuts thats $1.25 a lbs. with a pound of cashews at $1.80 a lbs. how many pounds of cashew do you need to make a $1.50 a lbs. mixture

Answer by jorel1380(3719)

12(1.25)+ x(1.80)=(12+x)1.50
15+1.80x=18+1.50x
.30x=3
x=3/.3=10 pounds of cashews need to be added to the mix to make a 22 lb. mixture worth $1.50 a pound..